OpenStudy (anonymous):

The equation of a line is x + 5y = 12. What is the y-intercept of the line?

OpenStudy (anonymous):

x+5y=12 -x -x 5y=-x+12(divide all this by 5) y= x/5 +12/5 The intercept would be 12/5